Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 16/5/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • TH chasing Thurlby-Thandar. Cause of failure is that there may still have been a few mA in the circuit when it was powered off with contactors still closed
      • The spare is installed and has been run into a short
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • Progress on Hall list List of Hall activities in the shutdown (Colin)
    • CWhyte to request survey of TOF1
  • DC update
  • MOM update
  • The tracker saga
  • DS pressure issue
  • Hydrogen system update
  • Run plan
    • Can't do much until we get the trackers operational (early next week?)
    • What we do then depends on the status of the lh2 system and result of hazop
    • should review run plan by end of week
  • AOB

Minutes

  • Actions
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • The spare is installed and has been run into a short
    • Hall Manager - water system change-over procedure (stands) - collar people on Friday
    • EOverton - Status of miceraid5 disk cleanup (stands) now has a script but needs to run it
    • Progress on Hall list List of Hall activities in the shutdown (Colin) DL were down on masse and made significant progress. Need to confirm if any remaining items
    • CWhyte to request survey of TOF1 - requested, next week as Tony off this week
  • New actions
    • SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings
    • AKurup to check which settings are being used by the active Alarm Handlers
    • SBoyd to conscript extra shift cover for the end of next week
  • DC update
    Activity has been predominantly the Tracker and the LH2, see below.
  • MOM update
    Last week did TOF calibration and Paolo's Q789-off TOF check.
    New Alarm Handler test stopped after miceecserv crash. Recently been getting a lot of beamline alarms, stemming from the glitches.
    "New" Alarm Handler has two aspects:
    • have new config files (from the spreadsheets) for the PVs
      in practice needs hand-coding details in the State Machine
    • have all-in-one UI
      has display issues - will leave until the configuration has been sorted out
  • The tracker cryostat water contamination débâcle
    Found ice in the cold trap coming out of the cryostat. Have got water out of cryo 3, moving to pumping out cryo 4.
    Cryo 1 also has a leak which is a separate issue.
    Will check the status of Cryo3 - i.e. dead channels - today. This will tell us if we're on the right track - if so ready next week.
    Have checked for and don't see an air leak in when pumping. Think the problem may have been a bad batch of helium at some point.
  • DS Shield Pressure saga
    The shield pressure has been falling slowly, but the rate of decline increased since powering. Trend appeared that we could have a problem before the end of the User Cycle.
    MCourthold has looked into this. We have experienced similar issues before. The drop is not as dramatic as in December. Belief is there is a small leak of air getting in. Concern is about T3125 which if it goes too high can release impurities from the cold trap inside the fridge.
    Not a concern on the 2-3 week timescale, but we should try cleaning the fridge gas in June to confirm we understand the issue.
  • Hydrogen system update
    Mods to second turret underway in R9 - spurious pipework removed, pressure measurement spigot by the end of the week. This turret to be installed ASAP in June for running LH2 in July.
    Cooling started last week in Hall: on Monday absorber at 21 K and first stage at 13 K. SW issue fixed. Ready for the Neon test once have staff from Thursday (after HAZOP).
    HAZOP ongoing, LOPA at end. Will determine if we can run LH2 at end of cycle.
  • Friday decision point
    Decide on Friday what to do for final 11 days. Should we conscript an extra shifter, to allow 24/7 running to start Thu/Fri? They may need to know this week to book accommodation, etc. and can be told there is uncertainty.
  • AOB
    All proposed work in the ISIS synch over the summer (June-Sept.) must be notified to Mike Krendler (via Henry) by the 19th May.

Updated by Nebrensky, Henry over 5 years ago · 17 revisions